It is desired that sufficient water be used to make 6 liters of alcohol 95% pure into a mixture that is 75% pure. How much water is needed
Answer by dkppathak(439)

It is desired that sufficient water be used to make 6 liters of alcohol 95% pure into a mixture that is 75% pure. How much water is needed
pure alcohol in 6 ltrs at 95%=6x95/100=5.7 leters
let K letters water is added in 95% solution
solution =6+x ltrs
as per given condition 75% of (6+x)=5.7
75/100 of (6+x)=5.7
.75 of (6+x)=5.7
6+x=5.7/.75 =7.6
x=7.6-6= 1.6 ltrs